# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ Abstract
Performance analysis of first-order algorithms with inexact oracle has gained recent attention due to various emerging applications in which obtaining exact gradients is impossible or computationally expensive. Previous research has demonstrated that the performance of accelerated first-order methods is more sensitive to gradient errors compared with non-accelerated ones. This paper investigates the nonasymptotic convergence bound of two accelerated methods with inexact gradients to solve deterministic smooth convex problems. Performance Estimation Problem (PEP) is used as the primary tool to analyze the convergence bounds of the underlying algorithms. By finding an analytical solution to PEP, we derive novel convergence bounds for Inexact Optimized Gradient Method (OGM) [39] and Inexact Fast Gradient Method (FGM) [16] with variable inexactness along iterations. Under the absolute error assumption, we derive bounds in which the accumulated errors are independent of the initial conditions and the trajectory of the sequences generated by the algorithms. Furthermore, we analyze the tradeoff between the convergence rate and accumulated error that guides finding the optimal stepsize. Finally, we determine the optimal strategy to set the gradient inexactness along iterations (if possible), ensuring that the accumulated error remains subordinate to the convergence rate.
@@ 1 Introduction
where $f$ is convex and has a Lipschitz continuous gradient. We assume the optimal value of (1) to be lower bounded, i.e., $f^{*}>-\infty$ with $f^{*}\triangleq\min_{\mathbf{x}}f(\mathbf{x})$ and $\mathcal{X}_{*}\triangleq\operatorname*{argmin}_{\mathbf{x}}f(\mathbf{x})$. We use $\mathcal{F}_{\mu,L}$ with $\mu\geq 0$ to denote the class of $\mu$ strongly convex functions with Lipschitz continuous gradient with constant $L$. This paper focuses on the class of merely convex functions with Lipschitz continuous gradient $f\in\mathcal{F}_{0,L}$. For this class of functions, the gradient descent (GD) method has the convergence rate $f(\mathbf{x}_{k})-f^{*}=\mathcal{O}(k^{-1})$. This rate can be improved to $\mathcal{O}(k^{-2})$ with Nesterovs’s fast gradient method (FGM) [51, 52, 50, 53]. Recently, using computer-aided analysis, the Optimized Gradient Method (OGM) has been proposed [22, 39] that improves the convergence rate of FGM by $\sqrt{2}$ constant.
The above rates are based on having access to the exact gradient at any point, which is not the case in many applications. The convergence rate deteriorates when there is an error in the gradient estimate. This paper considers gradient error in the absolute form as
@@
where the error $\mathbf{e}$ can be controlled by a parameter $\boldsymbol{\eta}\in\mathbb{R}^{r}$ through a positive function $h(\cdot)$ and the bound holds either deterministically, which is the focus of this work, or with high probability in stochastic settings. The parameter $\boldsymbol{\eta}$ represents, e.g., sampling size, computational effort, or accuracy of an approximation to achieve a desired gradient inexactness.
Below, the gradient oracle inexactness is illustrated through a few examples.
1.1 Examples of inexact gradient oracles
Gradient inexactness is generally the result of the oracle not being evaluated at the desired point. For instance, if some or all coordinates of the desired point are the solution of an optimization problem that is only solved to a suboptimal solution, the resulting gradient will be inexact, e.g., bilevel optimization. In some other scenarios, the desired point results from an expectation that can only be estimated resulting in gradient inexactness see, e.g., nested composition optimization.
Example 1.1 (Bilevel optimization [29]).
@@ 1.1 Examples of inexact gradient oracles
~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left\|\tilde{\mathbf{y}}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|,$
@@
Now let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations. From the convergence rate of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
~$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
Example 1.2 (Composition optimization).
Consider the nested composition optimization in the form of
Assume for any $\mathbf{v}$ and $\mathbf{x}$, $\left\|\nabla h(\mathbf{v})\right\|\leq C_{f}$ and $\left\|\nabla\mathbf{g}(\mathbf{x})\right\|\leq C_{g}$. Furthermore, assume $h(\mathbf{v})$ is Lipschitz smooth with constant $L_{f}$, $\mathbb{E}_{i}[\left\|\nabla\mathbf{g}_{i}(\mathbf{x})-\nabla\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\nabla}$, and $\mathbb{E}[\left\|\mathbf{g}_{i}(\mathbf{x})-\mathbf{g}(\mathbf{x})\right\|^{2}]\leq\sigma^{2}_{\mathbf{g}}$. Then with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
@@
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right).$
@@
Another common scenario is when $\nabla f$ is unattainable, but, they can be approximated by (inexact) function values, i.e., zero-order information, denoted by $\tilde{f}$. In many such scenarios, it can be assumed that $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq b_{f}$. When $f(\mathbf{x})$ has Lipschitz continuous gradient, the zero-order methods generate inexact gradient estimates with quantifiable error bounds (see [6] and references therein).
Example 1.3 (Gradient estimation via forward finite difference).
@@
Let $\vec{\mathbf{v}}_{i}\in\mathbb{R}^{d}$ denote the unit vector with the i-th element equal to 1, and $l>0$ be the finite difference interval. Define the i-th element of the gradient estimator as
[~$[\tilde{\nabla}f(\mathbf{x})]_{i}=\frac{\tilde{f}(\mathbf{x}+l\vec{\mathbf{v}}_{i})-\tilde{f}(\mathbf{x})}{l}.$
@@
The error of the gradient estimate is shown by [6] to be bounded as
~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\frac{\sqrt{d}Ll}{2}+\frac{2\sqrt{d}b_{f}}{l},$
@@
Another common approach for gradient approximation from (inexact) function values is based on Gaussian smoothing. This method randomly samples directions for gradient estimation; hence, the sample size can be some arbitrary number.
Example 1.4 (Gradient estimation via Gaussian smoothing).
Let $\{\mathbf{v}_{i}\}_{i=1}^{n}$ be i.i.d. random directions following the standard multivariate normal distribution, and define the gradient estimator as
~$\tilde{\nabla}f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{f}(\mathbf{x}+l\mathbf{v}_{i})-\tilde{f}(\mathbf{x})}{l}\mathbf{v}_{i}.$
@@
which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$. Furthermore,
~$\left\|\tilde{\nabla}f(\mathbf{x})-\mathbb{E}_{\{\mathbf{v}_{i}\}_{i=1}^{n}}[\tilde{\nabla}f(\mathbf{x})]\right\|^{2}\leq\frac{3d}{\epsilon n}\left(3\left\|\nabla f(\mathbf{x})\right\|^{2}+\frac{L^{2}l^{2}}{4}(d+2)(d+4)+\frac{4b_{f}^{2}}{l^{2}}\right)$
holds with probability at least $1-\frac{1}{\epsilon}$. In the later probabilistic bound, $h(\boldsymbol{\eta})$ is determined by $l$, $n$ and $b_{f}$ as $\boldsymbol{\eta}=(l,b_{f},n)$. This bound contains a constant, a relative error, and an absolute error terms.
1.2 Related works
The study on inexact gradient methods (without acceleration) dates back to 1987 [56], where two notions of inexactness were defined, i.e.,
~$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|$ $\leq\delta,\qquad$ (absolute error), (3)
@@ 1.2 Related works
Below, we review the literature in stochastic setups for non-accelerated methods and (in more detail) in deterministic setups for accelerated gradient-based methods.
Stochastic non-accelerated methods under biased oracles.
The above gradient inexactness assumptions have specifically been considered in stochastic setups where ~$\tilde{\nabla}f(\mathbf{x})$ is a biased stochastic estimate of $\nabla f(\mathbf{x})$. [1] studies the performance of the biased stochastic gradient descent (SGD) method under both absolute and relative bias. [36] considers setups where lower bias and variance require higher computational cost and investigates their tradeoff in stochastic optimization. Similarly, [35] analyzes bias-variance tradeoff in the convex bandit setting. [7] analyzes biased stochastic gradient-based methods under the absolute bias oracles with and without bias-variance tradeoff. [14] (and references therein) summarizes different bias assumptions and introduces a hybrid approach integrating absolute and relative biases. [47] considers biased gradient estimates with absolute error and proposes adaptive algorithms that adaptively set the bias throughout the iterations of the algorithms.
Accelerated gradient methods with inexact gradient oracles.
Nesterov-type acceleration techniques [51] with inexact gradient oracles have been analyzed under different inexact assumptions. Below, these assumptions and their connections with the absolute error (which is the focus of this paper) are reviewed; moreover, the convergence performance of deterministic accelerated methods with inexact gradients are analyzed under the corresponding assumptions.
An inexact generalization of Nesterov’s Fast Gradient Method (FGM), namely, iFGM, is presented in Algorithm 1 which obtains accelerated rates of convergence in convex setups. To simplify the exposition, we only consider the merely convex problems and we set the dual update step as gradient descent with projection.
Algorithm 1 Inexact Fast Gradient Method (iFGM)
0:  $\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $0<\alpha_{0}\leq 1$, $A_{0}=\alpha_{0}$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k}))$
3:     $\mathbf{z}_{k+1}=\Pi_{\mathcal{Q}}(\mathbf{x}_{0}-\sum_{i=0}^{k}\frac{\alpha_{i}}{L}\tilde{\nabla}f(\mathbf{x}_{i}))$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}^{2}\leq A_{k+1}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for

### v2

@@ Abstract
Performance analysis of first-order algorithms with inexact oracles has gained recent attention due to various emerging applications in which obtaining exact gradients is impossible or computationally expensive. Previous research has demonstrated that the performance of accelerated first-order methods is more sensitive to gradient errors compared with non-accelerated ones. This paper investigates the nonasymptotic convergence bound of two accelerated methods with inexact gradients to solve deterministic smooth convex problems. Performance Estimation Problem (PEP) is used as the primary tool to analyze the convergence bounds of the underlying algorithms. By finding an analytical solution to PEP, we derive novel convergence bounds of Generalized Optimized Gradient Method (GOGM) and Generalized Fast Gradient Method (GFGM) with inexact gradient oracles following the absolute error bound. The derived bounds allow varying oracle inexactness along the iterations; furthermore, their accumulated error terms are independent of the initial condition and any unknown parameters. Furthermore, we analyze the tradeoff between the vanishing term and the accumulated error in the convergence bound that guides finding the optimal stepsize. Finally, we determine the optimal strategy to set the gradient inexactness along iterations (if possible in a given application), ensuring that the accumulated error remains subordinate to the vanishing term.
@@ 1 Introduction
where $f$ is convex and has a Lipschitz continuous gradient. We assume the optimal value of (1) to be lower bounded, i.e., $f^{*}>$-$\infty$ with $f^{*}\triangleq\min_{\mathbf{x}}f(\mathbf{x})$ and $\mathcal{X}_{*}\triangleq\operatorname*{argmin}_{\mathbf{x}}f(\mathbf{x})$. We use $\mathcal{F}_{\mu,L}$ with $\mu\geq 0$ to denote the class of $\mu$ strongly convex functions with Lipschitz continuous gradient with constant $L$. This paper focuses on the class of merely convex functions with Lipschitz continuous gradient $f\in\mathcal{F}_{0,L}$. For this class of functions, the gradient descent (GD) method has the iteration complexity $f(\mathbf{x}_{k})-f^{*}=\mathcal{O}(k^{-1})$. This result can be improved to $\mathcal{O}(k^{-2})$ with Nesterov’s fast gradient method (FGM) [43, 44, 42, 45]. Recently, using computer-aided analysis, the Optimized Gradient Method (OGM) has been proposed [18, 32] that improves the complexity of FGM by $\sqrt{2}$ constant.
The above results are based on having access to the exact gradient at any point, which is not the case in many applications. The performance of the algorithms deteriorates when there is an error in the gradient estimate. We consider the gradient error as
@@
where the error $\mathbf{e}$ can be controlled by a parameter $\boldsymbol{\eta}\in\mathbb{R}^{r}$ through a positive function $h(\cdot)$ and the bound holds either deterministically, which is the focus of this work, or with high probability in stochastic settings. The parameter $\boldsymbol{\eta}$ represents, e.g., computational effort, the accuracy of an approximation, or sampling size to achieve a desired gradient accuracy.
1.1 Applications with inexact gradient oracles.
In some applications, gradient inexactness is the result of the oracle not being evaluated at the desired point. For instance, in bilevel optimization, one block of coordinates of the upper-level problem is the solution of the lower-level optimization problem. However, in practice, the lower-level problem can be solved to a suboptimal solution, which results in upper-level gradient inexactness. Similarly, in composition optimization, the desired point results from an expectation that in many scenarios can only be estimated, resulting in gradient inexactness. These two applications are discussed below.
Bilevel optimization.
@@
with $f$ being continuously differentiable and $g$ being two times continuously differentiable and strongly convex. To solve the problem with a gradient-based method, the gradient of the upper-level problem with respect to $\mathbf{x}$ is
$\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))=\nabla_{\mathbf{x}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{\mathbf{x}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{\mathbf{y}\mathbf{y}}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{\mathbf{y}}f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x})),$
@@
which requires solving the lower-level problem to optimality. When the lower-level problem is solved to a suboptimal point $\tilde{\mathbf{y}}(\mathbf{x})$ and the gradient is evaluated at this point instead of $\mathbf{y}^{*}(\mathbf{x})$, the resulting gradient is inexact. Let $\tilde{\mathbf{y}}_{k}(\mathbf{x})$ be the solution of the lower-level problem by the gradient descent method with stepsize $\frac{2}{\mu_{g}+L_{g}}$ after $k$ iterations. From the iteration complexity of GD for the class $\mathcal{F}_{\mu,L}$, the gradient inexactness can be bounded as
$\left\|\tilde{\nabla}f(\mathbf{x};\tilde{\mathbf{y}}_{k}(\mathbf{x}))-\nabla f(\mathbf{x};\mathbf{y}^{*}(\mathbf{x}))\right\|\leq C\left(\frac{Q_{g}-1}{Q_{g}+1}\right)^{k}\left\|\mathbf{y}_{0}-\mathbf{y}^{*}(\mathbf{x})\right\|,$
where $Q_{g}\triangleq L_{g}/\mu_{g}$. The above bound follows an exponential decay and represents the error bound in (2) with $\boldsymbol{\eta}$ being the iteration number $k$ – For more details see A.1.
Composition optimization.
Consider the nested composition optimization problem
$\min_{\mathbf{x}}\ f(\mathbf{x})\triangleq h(\mathbf{g}(\mathbf{x}))\quad\text{with}\quad\mathbf{g}(\mathbf{x})\triangleq\sum_{i=1}^{N}\mathbf{g}_{i}(\mathbf{x}),$
where $h:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\mathbf{g}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ are continuously differentiable, and $N$ is a large positive integer. From the chain rule, the exact gradient is $\nabla f(\mathbf{x})=\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))$. However, assume one can only approximate $\mathbf{g}$ and $\nabla\mathbf{g}$ through their minibach samples $\mathcal{S}_{\mathbf{g}}\subseteq[N]$ and $\mathcal{S}_{\nabla}\subseteq[N]$, respectively, where $N]\triangleq\{1,2,\cdots,N\}$. Then, under some Lipschitz continuity and bounded variance assumptions, as shown in Lemma A.1, with probability at least $1-\frac{1}{\epsilon}$, the inexact gradient estimate is bounded as
@@
$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{C_{1}}{\left|\mathcal{S}_{\nabla}\right|}+\frac{C_{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\right).$
This bound represents the error bound in (2) with $\boldsymbol{\eta}=\left(\left|\mathcal{S}_{\nabla}\right|,\left|\mathcal{S}_{\mathbf{g}}\right|\right)$. Furthermore, if $\mathcal{S}_{\nabla}=\mathcal{S}_{\mathbf{g}}$, for any fixed $\epsilon$, the bound follows a power law decay.
@@
Another scenario is when $\nabla f$ is unattainable, but it can be approximated by inexact function values (zero-order information), denoted by $\tilde{f}$. In many such scenarios, we can assume $\left|\tilde{f}(\mathbf{x})-f(\mathbf{x})\right|\leq b_{f}$. Below, we discuss a deterministic and stochastic zero-order method that generates inexact gradient estimates with quantifiable error bounds [4].
Gradient estimation via forward finite difference.
@@
Let $\mathbf{u}_{i}\in\mathbb{R}^{d}$ denote the unit vector with the $i$-th element equal to 1, and $l>0$ be the finite difference interval. Define the $i$-th element of the gradient estimator as
$\tilde{\nabla}f(\mathbf{x})]_{i}=\frac{\tilde{f}(\mathbf{x}+l\mathbf{u}_{i})-\tilde{f}(\mathbf{x})}{l}.$
@@
When $f(\mathbf{x})$ has Lipschitz continuous gradient, the error of the gradient estimate is shown by [4] to be bounded as
$\left\|\tilde{\nabla}f(\mathbf{x})-\nabla f(\mathbf{x})\right\|\leq\frac{\sqrt{d}Ll}{2}+\frac{2\sqrt{d}b_{f}}{l},$
@@
Gradient estimation via Gaussian smoothing.
Let $\{\mathbf{v}_{i}\}_{i=1}^{n}$ be i.i.d. random directions following the standard multivariate normal distribution, and define
$\tilde{\nabla}f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{f}(\mathbf{x}+l\mathbf{v}_{i})-\tilde{f}(\mathbf{x})}{l}\mathbf{v}_{i}.$
@@
which represents the error bound (2) with $\boldsymbol{\eta}=(l,b_{f})$.
1.2 Effect of oracle inexactness on acceleration
A careful review of the literature shows that the effect of oracle inexactness on acceleration is addressed through 1) development of a specific error condition, and 2) analysis of a common (or customized) accelerated method under that error condition with (or in a few cases without) extra assumption. A detailed literature review on accelerated methods under different gradient error conditions is relegated to Section 1.4.
This paper considers gradient inexactness in the form of the absolute error defined as [48]
@@ 1.2 Effect of oracle inexactness on acceleration
Motivated by the application discussed in Section 1.1, (AE) can be considered as one of the most common forms to quantify the gradient error. Furthermore, as discussed in Section 1.4, obtaining other error conditions from (AE), e.g., BIE, IFO, etc., requires extra assumptions that limit their applications.
Nesterov’s Fast Gradient Method (FGM) and its variations have been extensively studied under various inexact gradient assumptions, which we refer to as iFGM. The standard FGM update generates sequences $\{\mathbf{y}_{k}\}$ and $\{\mathbf{z}_{k}\}$, with the gradient evaluated at a convex combination of these points. The detailed update rule is
$\begin{cases}\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k}),\\ \mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{1}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k}),\\ \mathbf{x}_{k+1}=(1-\frac{1}{\alpha_{k+1}})\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},\end{cases}$
where $\alpha_{0}=1$, $\alpha_{k+1}=\frac{1+\sqrt{1+4\alpha_{k}^{2}}}{2}$, and $\tilde{\nabla}$ denotes the inexact gradient. When the exact gradient is available, FGM achieves the optimal convergence rate of $\mathcal{O}(K^{-2})$. However, when the gradient is inexact, this fast convergence also accelerates error accumulation, undermining the method’s performance.
To address this, a generalized version of iFGM (iGFGM, Algorithm 1) is often considered. This approach controls the increase of $\alpha_{k}$ to obtain a better tradeoff between the convergence rate and the accumulated error. For instance, under a specific assumption of the gradient inexactness (see details in Section 1.4), Devolder et al. [12] established the convergence result of iGFGM as
$\mathcal{O}\left(\frac{1}{A_{K}}+\frac{\sum_{i=0}^{K}A_{i}\delta_{\mathbf{x}_{i}}}{A_{K}}\right),$
where $\delta_{x_{i}}$ represents the gradient error at $x_{i}$ and $A_{i}=\sum_{j=1}^{i}\alpha_{j}$ with $\alpha_{i}^{2}\leq A_{i}$.
While iGFGM has been widely investigated under error conditions other than (AE) (except for one work [58] which is discussed in Section 1.4), the Optimized Gradient Method (OGM) with an inexact gradient, i.e., iOGM, has received less attention. A similar tradeoff between convergence rate and accumulated error exists for iOGM. In this paper, we consider the generalized version [35], which we term iGOGM (Algorithm 2).
We note that the difference between the iGFGM and iGOGM methods is in their step 3, where iGOGM’s stepsize is two times larger. When $\lambda_{k}=1$, the $\mathbf{x}_{k+1}$ updates (step 5) in both iGFGM and iGOGM are simplified as $\mathbf{x}_{k+1}=\left(1-\frac{1}{\alpha_{k+1}}\right)\mathbf{y}_{k+1}+\frac{1}{\alpha_{k+1}}\mathbf{z}_{k+1},$ and algorithms reduce to iFGM and iOGM, respectively.
[…]
1.3 Contributions
In this paper, we analyze nonasymptotic convergence bounds of two accelerated gradient methods, namely Generalized Fast Gradient Method (iGFGM, Algorithm 1) and Generalized Optimized Gradient Method (iGOGM, Algorithm 2), under inexact gradient oracles satisfying the absolute error condition (AE). The (AE) error bound finds a range of applications, a few of which are discussed in Section 1.1. These bounds are established through the Performance Estimation Problem (PEP) technique and, unlike previous works, are independent of unknown quantities–see the first bullet below for more details. Furthermore, our analysis allows variable error along the iterations of the algorithms; hence, the established bounds allow exploiting the tradeoff between the per-iteration cost to control the bias and the total cost to obtain the optimal oracle inexactness schedule. Our key contributions can be summarized as follows:
  • As summarized in Table 1, existing methods are limited by their reliance on strong assumptions (e.g., BIE, IFO, IFO-q, and RE) which are generally not easily verifiable in different applications. Furthermore, some previous analyses depend on unquantifiable terms, e.g., the dependence of the bound in [58] on unquantifiable $\tilde{R}_{K}$, the radius of the set containing the algorithm’s trajectory. Hence, it is not possible to evaluate the accumulated error in [58] and to determine the algorithm’s optimal parameter settings. Our work closes this gap by deriving a quantifiable expression for the accumulated error under the AE condition.
  • We provide the convergence bound of iGOGM and iGFGM under the absolute error (AE) condition in Section 2.1. The derived convergence bounds consist of two components: the diminishing component independent of the oracle’s error and the component containing the accumulated error. Interestingly, the accumulated error is independent of the initial condition $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|$, and is determined solely by the Lipschitz constant and the stepsize. This result advances the convergence bound of [58], presented in (4), in the sense that it eliminates the dependence on the unknown parameter $\tilde{R}_{k}$. Furthermore, unlike Nabou et al. [41]’s inexactness condition for the gradient to be a subgradient, we do not require such a condition, and we yet achieve comparable convergence guarantees.
  • In Section 2.2, we analyze the established convergence bounds from two different perspectives: i) we investigate the tradeoff between the vanishing term and the accumulated error; ii) we establish the optimal inexactness schedule, i.e., the minimal total cost to control the error, while preserving the accelerated convergence bound. The proofs of our main convergence results for iGOGM and iGFGM are presented in Sections 4 and (A.2), respectively.
  • We derive an analytical feasible solution to the dual of the relaxed semidefinite programming formulation of the Performance Estimation Problem (PEP) for first-order algorithms with inexact (AE) oracles in Section 4. The approach used to find this solution inspires our theoretical proof of the convergence bound. Notably, this proof can be understood without prior knowledge of the PEP technique and could be of independent interest.
The codes for our numerical experiments as well as those to verify the algebraic steps in our proofs are available at https://github.com/Yin-LIU/Inexact-Acceleration-PEP.
Notations
Vectors and matrices are denoted by bold and capitalized bold letters, respectively. Sets are denoted by calligraphic letters. In the absence of additional instruction, a bold letter coupled with an arrow, e.g., $\vec{\mathbf{u}}$, generally denotes a standard basis vector, which has exactly one element equal to 1 while other elements are 0. $\mathbf{M}_{i,j}$ stands for the $(i,j)$-th element of the matrix $\mathbf{M}$. The notation $\mathbf{M}^{\backslash[i]}$ refers to the submatrix of $\mathbf{M}$ obtained by removing its $i$-th column and row. $\mathbf{M}]_{n\times m}$ indicates the shape of the matrix and $\mathbf{M}]$ is used to emphasize the item inside the brackets is a matrix. $\left\|\cdot\right\|$ denotes the $l_{2}$ norm for vectors and Frobenius norm for matrices. $\log(\cdot)$ represents the natural logarithm.
1.4 Related work
Below, we discuss the literature related to accelerated methods under various oracle inexactness conditions.